# Improving Implicit Regularization of SGD with Preconditioning for Least Square Problems

### v1

@@ 4.5 Empirical Study
We perform experiments on Gaussian least square problem. [^1] Anonymous link to the implementation:
https://anonymous.4open.science/r/presgd-4778
We consider $6$ problem instances, which are the combinations of $2$ different covariance matrices $\mathbf{H}$, $\lambda_{i}=i^{-1}$ and $\lambda_{i}=i^{-2}$, and $3$ different ground-truth vectors $\mathbf{w}^{*}$, $\mathbf{w}^{*}[i]=1$, $\mathbf{w}^{*}[i]=i^{-1}$, and $\mathbf{w}^{*}[i]=i^{-10}$. This empirical study aims to answer the following key questions:
@@
    Can SGD with preconditioning indeed outperform ridge?
@@
    Does the benefit of preconditioning persist when the preconditioning matrix is empirically estimated?
We compare the generalization performance of four optimization methods: SGD, Ridge Regression, Preconditioned-SGD (PreSGD), and PreSGD with estimation (PreSGD-Est) across the 6 problem instances. A set of unlabelled data of the same size as the training set is drawn from the same problem instance to estimate $\widetilde{\mathbf{G}}$ in PreSGD-Est. The results of these comparisons are given in Figure 1. In cases where Ridge Regression consistently outperforms SGD, such as when $\mathbf{w}^{*}[i]=1$, both PreSGD and PreSGD-Est demonstrate the ability to achieve comparable performance to Ridge Regression. Note that when $\mathbf{w}^{*}[i]=1$, the bias error will be large. Therefore the improvement of PreSGD and PreSGD-Est in this case well supports our intuition and theory regarding the design of preconditioning. Conversely, in scenarios where SGD exhibits comparable or superior performance to ridge regression, such as when $\mathbf{w}^{*}[i]=i^{-1}$ or $\mathbf{w}^{*}[i]=i^{-10}$, both PreSGD and PreSGD-Est manage to further enhance the generalization performance. These empirical findings support the hypotheses posed in our theoretical analysis, validating the effectiveness of preconditioning in improving the generalization performance of SGD.
5 Conclusion
This paper conducts in-depth exploration of implicit regularization in SGD with preconditioning. Our analysis extends the boundaries of existing understanding by characterizing the excessive risk associated with both SGD and ridge regression within the context of preconditioning. We show that preconditioned SGD consistently outperforms both ridge regression with and without preconditioning, effectively closing the performance gap between SGD and ridge. Furthermore, we illustrate that our proposed preconditioned matrix can be robustly estimated using readily available, inexpensive, and unlabeled data, affirming its practical feasibility and real-world applicability.
In essence, our findings establish that SGD with preconditioning consistently achieves better generalization performance compared to ridge regression. This research underscores the vital role of SGD in machine learning and emphasizes its potential for further improvement through the utilization of preconditioning techniques. Moving forward, our future research will explore whether preconditioning can also enhance the implicit regularization of SGD in the context of other linear and potentially nonlinear models.
@@ Appendix C Analysis of preconditioned SGD with preconditioned ridge regression
In the previous appendix, we have proved that SGD with preconditioning can indeed outperform the standard ridge regression. In this section, we compare the excessive risk of preconditioned SGD and preconditioned ridge. We aim to show that for a given precondition matrix $\mathbf{M}$, our proposed design of preconditioned matrix $\mathbf{G}$ for SGD with slight modification can outperform the excessive risk of precondition ridge.
@@ Appendix D Analysis of precondition SGD with empirical estimation
In the previous appendix, we have shown that SGD with preconditioning can consistently outperform ridge regression with and without preconditioning. In the previous proof, we assume access to the precise information of the data covariance matrix $\mathbf{H}$.
@@
$\leq\sum_{i}\frac{\lambda_{i}}{1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}$
@@
$\frac{\lambda_{i}}{1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}$ $=\frac{\lambda_{i}}{1+\frac{C\lambda_{i}}{1+\beta\sum_{j>\widehat{k}}\lambda_{j}}}$
$\leq\frac{\lambda_{i}}{1+\frac{C\lambda_{i}}{\lambda_{i}}}$
@@
$\frac{\lambda_{i}}{1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}$ $=\frac{\lambda_{i}}{\frac{1+\beta\sum_{j>\widehat{k}}\lambda_{j}+C\lambda_{i}}{1+\beta\sum_{j>\widehat{k}}\lambda_{j}}}$
@@
$\frac{\lambda_{i}}{1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}$ $\leq\frac{\lambda_{i}\lambda_{i}}{\lambda_{i}+C\lambda_{i}}$

### v3

@@ 4.5 Empirical Study
We perform experiments on Gaussian least square problem [^4] Anonymous link to the implementation: https://github.com/jwsu825/presgd. We consider $6$ problem instances, which are the combinations of $2$ different covariance matrices $\mathbf{H}$, $\lambda_{i}=i^{-1}$ and $\lambda_{i}=i^{-2}$, and $3$ different ground-truth vectors $\mathbf{w}^{*}$, $\mathbf{w}^{*}[i]=1$, $\mathbf{w}^{*}[i]=i^{-1}$, and $\mathbf{w}^{*}[i]=i^{-10}$. This empirical study aims to answer the following key questions:
@@
    Can SGD with preconditioning indeed be consistently comparable with ridge?
@@
    Is preconditioning still beneficial when the preconditioning matrix is empirically estimated?
We compare the generalization performance of four optimization methods: SGD, Ridge Regression, Preconditioned-SGD (PreSGD), and PreSGD with estimation (PreSGD-Est) across the 6 problem instances. A set of unlabelled data of the same size as the training set is drawn from the same problem instance to estimate $\widetilde{\mathbf{G}}$ in PreSGD-Est. The results of these comparisons are given in Figure 1. In cases where Ridge Regression outperforms SGD, such as when $\mathbf{w}^{*}[i]=1$, both PreSGD and PreSGD-Est demonstrate the ability to achieve comparable performance to Ridge Regression. Note that when $\mathbf{w}^{*}[i]=1$, the bias error will be large. Therefore the improvement of PreSGD and PreSGD-Est in this case well supports our intuition and theory regarding the design of preconditioning. Conversely, in scenarios where SGD exhibits comparable or superior performance to ridge regression, such as when $\mathbf{w}^{*}[i]=i^{-1}$ or $\mathbf{w}^{*}[i]=i^{-10}$, both PreSGD and PreSGD-Est manage to further enhance the generalization performance. These empirical findings support the hypotheses posed in our theoretical analysis, validating the effectiveness of preconditioning in improving the generalization performance of SGD.
5 Discussion
Conclusion. We conduct an in-depth exploration of implicit regularization in SGD with preconditioning. Our analysis extends the boundaries of existing understanding by characterizing the excessive risk associated with both SGD and ridge regression within the context of preconditioning. We show that preconditioned SGD can be consistently comparable with ridge regression with and without preconditioning, effectively closing the performance gap between SGD and ridge. Furthermore, we illustrate that our proposed preconditioned matrix can be robustly estimated using readily available, inexpensive, and unlabeled data, affirming its practical feasibility and real-world applicability. This research underscores the vital role of SGD in machine learning and emphasizes its potential for further improvement through the utilization of preconditioning techniques.
Limitation and Future Work. Our primary objective in this study was to address an open question from the literature, and we consequently focused on the linear regression problem to maintain consistency with the setting of the previous study. Looking ahead, a natural progression for future research involves expanding our investigation beyond linear models and SGD. We aim to explore whether preconditioning can also enhance the implicit regularization of SGD or similar algorithms within the context of other linear and potentially nonlinear models. More specifically, we are considering the following areas for future research: 1) extending the analysis to explore preconditioned SGD for locally quadratic models, where decaying step sizes are essential for reducing risk to zero; 2) examining the application of preconditioning to SGD in the context of matrix regression (Wu et al., 2023a); and 3) analyzing the effects of preconditioning on SGD and GLM-tron (Kakade et al., 2011) in ReLU regression (we have presented preliminary results and discussions for this direction in Appendix E.2).
@@ Appendix C Analysis of preconditioned SGD with preconditioned ridge regression
In the previous appendix, we have proved that SGD with preconditioning can indeed match the performance of the standard ridge regression. In this section, we compare the excessive risk of preconditioned SGD and preconditioned ridge. We aim to show that for a given precondition matrix $\mathbf{M}$, our proposed design of preconditioned matrix $\mathbf{G}$ for SGD with slight modification can remain comparable to the excessive risk of precondition ridge.
@@ Appendix D Analysis of precondition SGD with empirical estimation
In the previous appendix, we have shown that SGD with preconditioning can consistently match ridge regression with and without preconditioning. In the previous proof, we assume access to the precise information of the data covariance matrix $\mathbf{H}$.
@@
$\mathbf{K}$ $=\mathbf{H}^{1/2}\widetilde{\mathbf{G}}\mathbf{H}^{1/2}$
$=\mathbf{H}^{1/2}(\beta\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\top}+\mathbf{I})^{-1}\mathbf{H}^{1/2}$
$=(\beta\mathbf{H}^{-1/2}\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\top}\mathbf{H}^{-1/2}+\mathbf{H}^{-1})^{-1}.$
@@
$\leq\sum_{i}\lambda_{i}/\left({1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}\right)$
@@
$\lambda_{i}/\left({1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}\right)$ $=\lambda_{i}/\left({1+\frac{C\lambda_{i}}{1+\beta\sum_{j>\widehat{k}}\lambda_{j}}}\right)$
$\leq\lambda_{i}\left({1+\frac{C\lambda_{i}}{\lambda_{i}}}\right)$
@@
$\lambda_{i}/\left({1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}\right)$ $=\lambda_{i}/\left(\frac{1+\beta\sum_{j>\widehat{k}}\lambda_{j}+C\lambda_{i}}{1+\beta\sum_{j>\widehat{k}}\lambda_{j}}\right)$
@@
$\lambda_{i}/\left({1+C\beta\lambda_{i}\cdot\frac{1}{\beta+\beta^{2}\sum_{j>\widehat{k}}\lambda_{j}}}\right)$ $\leq\frac{\lambda_{i}\lambda_{i}}{\lambda_{i}+C\lambda_{i}}$